Do not use thousand separator, use 2 decimal places only (i.e. 1000.00 instead of 1,000).

What is the SYD five-year depreciation schedule for a truck that was purchased for Php4M with a lifetime of 5 years and has a residual value of Php750,000?

Year 1 2 3 4 5
SYD Fraction (in %) % % % % %
Depreciation
Answer #1

Cost of truck = 4M or 4 million

= 4*1000000

= Php 4000000

Residual Value = Php 750000

Useful life = 5 Years

Step -1 Calculation of the sum of years digit in an assets useful life -

For an assets having useful life of 5 Years, the sum of years digit will be calculated as follows:

Sum of years digit = 5+4+3+2+1

= 15

Step - 2 Calculation of depreciable amount

Depreciable amount = Assets acquisition cost - Residual value

= 4000000-750000

= 3250000

Step -3 Calculation of the un-depreciated useful life

year 1 year 2 year 3 year 4 year 5
Undepreciated Useful life 5 4 3 2 1

Step 4 Calculation of the Depreciation expense -

Depreciation expense = Undepreciated useful life at the end of each year/ sum of years digit*Depreciable Amount

Year 1 = 5/15*3250000

= 1083333.33

Year 2 = 4/15*3250000

= 866666.67

Year 3 = 3/15*3250000

= 650000

Year 4 = 2/15*3250000

= 433333.33

Year 5 = 1/15*3250000

= 216666.67

Year 1 Year 2 Year 3 Year 4 Year 5
a Undepreciated Useful life 5 4 3 2 1
b sum of years digit 15 15 15 15 15
c SYD Fraction (%) (a/b) 33.33% 26.67% 20.00% 13.33% 6.67%
d Depreciable base (acquisition cost - residual value) 3250000 3250000 3250000 3250000 3250000
e Depreciation Php (c*d) 1083333.33 866666.67 650000.00 433333.33 216666.67

Please check with your answer and let me know.